# Unlearning as multi-task optimization: A normalized gradient difference approach with an adaptive learning rate

### v2

@@ 1 Introduction
Inspired by the simplicity and strong empirical performance of linear scalarization methods,222xin2022current [51] demonstrate that linear scalarization outperforms, or is at least on par with, other MTO approaches across various language and vision experiments which minimize a linearly weighted average of task losses, we propose an LLM unlearning method, NGDiff, based on dynamic scalarization, and analyze its theoretical properties. Building on the analysis, we introduce an automatic learning rate adaptation method tailored for LLM unlearning.
@@ 2 Related Work
Multi-task optimization In NLP, multi-task learning [56] typically refers to building a model that can perform well on multiple tasks simultaneously by sharing representations, introducing constraints, or combining multiple learning objectives [^3] This often involves optimizing a form of static linear scalarization, as introduced in the next section. Multi-task optimization, on the other hand, focuses on a slightly different concept – optimizing two distinct learning objectives simultaneously. The key challenge is how to balance the trade-off among objectives during the optimization procedure by modifying the per-task gradients (e.g. PCGrad [52], RLW [30], IMTL [33]). Several recent works have studied Pareto frontier and optimally in context of NLP tasks (e.g., multi-lingual machine translation [8, 51] and NLP fairness [19]). While optimization is often treated as a black-box tool in NLP research, studying optimization provides deeper insights and inspires new algorithms. [^4] An example is the Baum-Welch algorithm, originally proposed to estimate the parameters of HMMs and applied in speech recognition in 1970s before it was recognized as an instance of the EM algorithm [15]. It was later identified as a special case of a broader class of convex-concave optimization (CCCP) [53]. This connection inspired new designs, such as the unified EM [42].
@@ 3.1 Static linear scalarization
Figure 2: Gradient space in 2-dimension. $\color[rgb]{1,0,0}\bm{g}_{\textup{F}}$ is the forgetting gradient and $\color[rgb]{0,0,1}\bm{g}_{\textup{R}}$ is the retaining gradient, each with a perpendicular dashed line. Yellow area is the linear span (Eq. (3)) by scalarization. Green area is positively correlated to $\bm{g}_{\textup{R}}$ and negatively correlated to $\bm{g}_{\textup{F}}$ by Eq. (7), whereas NGDiff always stays within this green area at each iteration.
@@ 4.2 Normalized gradient difference
To satisfy Eq. (7) , we propose a normalized gradient difference method (NGDiff) to dynamically set $c_{t}=\frac{1/\|\bm{g}_{\textup{R}}\|}{1/\|\bm{g}_{\textup{R}}\|+1/\|\bm{g}_{\textup{F}}\|}\Longrightarrow\bm{g}_{\textup{NGDiff}}(\bm{g}_{\textup{R}},\bm{g}_{\textup{F}}):=\frac{\bm{g}_{\textup{R}}}{\|\bm{g}_{\textup{R}}\|}-\frac{\bm{g}_{\textup{F}}}{\|\bm{g}_{\textup{F}}\|}.$ In words, we normalize the retaining and forgetting gradients [^6] We illustrate in Appendix B that NGDiff is critically different and simpler than GradNorm.[10].
We analyze NGDiff as follows. First, we show the condition in Eq. (7) is satisfied at all iterations in the following lemma (proof in Appendix G):
@@ 5.1 Setup
Unlearning methods We compare NGDiff with 4 baselines. The first baseline method is the target model without any unlearning, while the remaining three are popular unlearning methods.
@@ F.3 Hyper-parameter Settings
To finetune a targeted model with the full dataset, we use the optimizer Adam with a learning rate of $\eta=\{10^{-5},2*10^{-5}\}$, a training batch size of {16, 32}, and train 25 epochs for all language models. For the unlearning process, we use the optimizer Adam with a learning rate $\eta=\{10^{-5},5*10^{-5},10^{-4}\}$, and train 15 epochs for all unlearning methods.

### v3

@@ 1 Introduction
Inspired by the simplicity and strong empirical performance of linear scalarization methods,222[51] demonstrate that linear scalarization outperforms, or is at least on par with, other MTO approaches across various language and vision experiments which minimize a linearly weighted average of task losses, we propose an LLM unlearning method, NGDiff, based on dynamic scalarization, and analyze its theoretical properties. Building on the analysis, we introduce an automatic learning rate adaptation method tailored for LLM unlearning.
@@ 2 Related Work
Multi-task optimization In NLP, multi-task learning [56] typically refers to building a model that can perform well on multiple tasks simultaneously by sharing representations, introducing constraints, or combining multiple learning objectives. This often involves optimizing a form of static linear scalarization, as introduced in the next section. Multi-task optimization, on the other hand, focuses on a slightly different concept – optimizing two distinct learning objectives simultaneously. The key challenge is how to balance the trade-off among objectives during the optimization procedure by modifying the per-task gradients (e.g. PCGrad [52], RLW [30], IMTL [33]). Several recent works have studied Pareto frontier and optimally in context of NLP tasks (e.g., multi-lingual machine translation [8, 51] and NLP fairness [18]). While optimization is often treated as a black-box tool in NLP research, studying optimization provides deeper insights and inspires new algorithms. [^3] An example is the Baum-Welch algorithm, originally proposed to estimate the parameters of HMMs and applied in speech recognition in 1970s before it was recognized as an instance of the EM algorithm [14]. It was later identified as a special case of a broader class of convex-concave optimization (CCCP) [53]. This connection inspired new designs, such as the unified EM [42].
@@ 3.1 Static linear scalarization
Figure 2: Gradient space in 2-dimension. $\color[rgb]{1,0,0}\bm{g}_{\textup{F}}$ is the forgetting gradient and $\color[rgb]{0,0,1}\bm{g}_{\textup{R}}$ is the retaining gradient, each with a perpendicular dashed line. Yellow area is the linear span (Eq. (3)) by scalarization. Green area is positively correlated to $\bm{g}_{\textup{R}}$ and negatively correlated to $\bm{g}_{\textup{F}}$ by Eq. (7), whereas NGDiff always stays within this green area at each iteration by Lemma 4.
@@ 4.2 Normalized gradient difference
To satisfy Eq. (7) , we propose a normalized gradient difference method (NGDiff) to dynamically set $c_{t}=\frac{1/\|\bm{g}_{\textup{R}}\|}{1/\|\bm{g}_{\textup{R}}\|+1/\|\bm{g}_{\textup{F}}\|}\Longrightarrow\bm{g}_{\textup{NGDiff}}(\bm{g}_{\textup{R}},\bm{g}_{\textup{F}}):=\frac{\bm{g}_{\textup{R}}}{\|\bm{g}_{\textup{R}}\|}-\frac{\bm{g}_{\textup{F}}}{\|\bm{g}_{\textup{F}}\|}.$ In words, we normalize the retaining and forgetting gradients [^5] We illustrate in Appendix B that NGDiff is critically different and simpler than GradNorm.[10].. We analyze NGDiff as follows. First, we show the condition in Eq. (7) is satisfied at all iterations in the following lemma (proof in Appendix G):
@@
$\bm{g}_{\textup{NGDiff}}(\bm{g}_{\textup{R}},\bm{g}_{\textup{F}})$ satisfies Eq. (7) for any $\bm{g}_{\textup{R}}\in\mathbb{R}^{d}$ and $\bm{g}_{\textup{F}}\in\mathbb{R}^{d}$. In words, $\bm{g}_{\textup{NGDiff}}$ is always positively correlated to $\bm{g}_{\textup{R}}$ and negatively correlated to $\bm{g}_{\textup{F}}$.
@@ 5.1 Setup
Unlearning methods We compare NGDiff with 4 baselines. The first baseline method is the target model without any unlearning, while the remaining three are the state-of-the-art unlearning methods.
@@ 6 Conclusion and Discussion
Acknowledgments
We would like to thank Shankar Ananthakrishnan, Fabian Triefenbach and Jianhua Lu from Amazon AGI Foundations team for providing feedback on this paper and supporting this work.
@@ F.3 Hyper-parameter Settings
To finetune a targeted model with the full dataset, we use the optimizer Adam with a learning rate of $\eta=\{10^{-5},2*10^{-5}\}$, a training batch size of {16, 32}, and train 25 epochs for all language models. For the unlearning process, we use the optimizer Adam with a learning rate $\eta=\{10^{-5},5*10^{-5},10^{-4}\}$, and train 15 epochs for all unlearning methods. For the Phi-1.5, Falcon-1B, and GPT2-XL models, we perform full-model parameter tuning. For the Llama2-7B and Mistral-7B models, we apply the LoRA training method [20] with rank = 8.
@@ Appendix G Proofs
Remark G.1.
Proof of Remark G.1.
We extend our original complexity analysis in Remark 4.2 and provide quantitative analysis of computation overheads in FLOPs (floating point operations). Specifically, assume it takes around N FLOPS to perform one forward pass on one example and 2N FLOPS to back-propagate. The basic GDiff requires roughly 6BTN FLOPS to run with batch size B and total number of iterations T, because it needs 1 forward and 1 backward for the retain set and another for the forget set. Our learning-rate-free NGDiff requires about (6BTN+4BTN/10), hence a 6.6% increase to unlearn the data. We assure that the extra forward passes do not add memory burden, because they are in gradient-free mode (e.g. under torch.no_grad() mode). All in all, NGDiff is almost as efficient as GDiff and other unlearning methods, as the target unlearning corpus is usually relatively small. This has been empirically observed by our large-scale (up to 7B) model unlearning experiments. ∎